# 2Leibniz-Institut für Astrophysik Potsdam (AIP), An der Sternwarte 16, 14482 Potsdam, Germany

### v1

@@ 4.3.3 CV sub-types and potential selection bias
Thirdly we considered the BOSS QA sample. This is essentially an unseen sample and shows very little bias. The overall recovery rate for the BOSS QA sample is 91 per cent.
@@
Disc Sp P
J000155.14–670743.3 EF Tuc 4707482485122305536 3.5: 14.85 $1329$ NL 26 18 15
J000600.15+012129.8 2738755406045571968 1.52(10) 19.57 $352$ WZ Sge: 23 23 23
J000633.70–690033.8 HV 8001 4706297108508032128 1.9 15.1 $241$ SU UMa 5 10 10
J000844.33–014014.8 Gaia15abi 2544817841421994624 18.15 $1236$ NL:/Polar 30 23
J001204.50+020129.6 2546760133008036096 3.281 16.45 $1667$ NL 12 3 1
J002243.55+061002.7 MGAB-V295 2747622177049417344 1.9(1): 19.61 $2113$ SU UMa 19 7 7
J002257.65+614107.5 V1033 Cas 430446643235641472 4.033 16.58 $1557$ IP 16 25 25
J002728.02–010828.7 EN Cet 2541910801397761152 1.424 20.76 $1218$ WZ Sge 2 27 28
J002848.83+591722.0 V0709 Cas 428101969045877504 5.3329 14.39 $725$ IP 24 24 17
@@
J014227.07+001729.8 2510205490257050496 1.88(4) 19.89 $697$ Polar: 14 14
This is a small sample of the data. The full table can be found in the supplementary data. References:  1 This work, 2 Bao Supernova Survey et al. (1997), 3 Luo et al. (2018), 4 Kato et al. (2015), 5 Luyten (1933), 6 Stefanov & Stefanov (2023), 7 Inight et al. (2023a), 8 Drake et al. (2009b), 9 Szkody et al. (2018), 10 Augusteijn et al. (2010), 11 Southworth et al. (2010), 12 vsnet-chat 7946, 13 AAVSO, 14 Carbon et al. (2017), 15 CBA512, 16 Halpern & Mirabal (2006), 17 Thorstensen et al. (2010), 18 Stobie et al. (1995), 19 MGAB Variable Star Catalog, 20 Szkody et al. (2014), 21 Southworth et al. (2008), 22 Shappee et al. (2014), 23 Inight et al. (2023b), 24 Motch et al. (1996), 25 Bonnet-Bidaud et al. (2007), 26 Chen et al. (2001), 27 Szkody et al. (2005), 28 Dillon et al. (2008), 29 Szkody et al. (2004), 30 Hodgkin et al. (2021a), 31 VSNET 12318, 32 Southworth et al. (2007)
@@ 5 CVs found in SDSS-V
505  CVs, consisting of the 478  CVs identified by the CNN plus the 27  identified manually were further analysed following the methods in Inight et al. (2023a, b). This analysis of the SDSS-Vspectroscopy and ancillary data of the 505  CVs in our sample yielded 62  new CVs, spectroscopic confirmation of 243 candidates and 68  new or improved orbital periods. Prior to SDSS-V, one of the authors (JRT) obtained time-resolved spectroscopy using the MDM observatory at Kitt Peak (see Thorstensen 2020 for more details on the instrument and data reduction techniques) of a number of our CVs. We present here the so far unpublished periods resulting from these observations. The full set of data is included in the supplementary data and a sample is shown in Table LABEL:tab:cvlist. Whilst analysing the 505  CVs in SDSS-V some CVs with unusual or noteworthy characteristics were identified and these are discussed below. For simplicity the names of CVs have been abbreviated in the text by dropping the arcseconds; so for example SDSS J015543.47+002807.4 becomes J0155+0028.
@@
The CRTS light curve (Fig. LABEL:fig:J0155+0028) captured the system mainly in a high state ($16-18$ mag), in contrast, the more recent ZTF data shows the system predominantly in a low state ($\simeq 20$ mag), with two relatively short high states. The transitions between high and low states can occur on time scales of days .
@@
This system is also known as ATO J119.2035-12.7816. The SDSS spectrum shows double-peaked Balmer and He i emission lines. The object is on the main sequence in the HR diagram just below the region occupied by G-type stars. GALEX observations show that the system is fairly bright in the ultraviolet. The ZTF light curve exhibits outbursts. We obtained a period of $10.79(4)$ h from the TESS light curve. This appears to be an U Gem dwarf nova, possibly with an evolved donor although an unevolved G-type donor would also be consistent.
@@
This system is also known as BMAM-V636. The SDSS spectrum shows Balmer, He i and He ii emission lines. J1618+0309 is a known polar (BMAM-V636). Cyclotron humps are evident in the spectrum. (Fig. LABEL:fig:J1618+0309) indicating $B\simeq 37$ MG.
@@
This is the classical nova V728 Sco, which erupted in 1862. The SDSS spectrum shows double-peaked Balmer, He i and He ii emission lines. J1739–4527 is located close to the main sequence in the HR diagram. There is no light curve data available. We concur with Tappert et al. (2012) that J1739–4527 now exhibits the characteristics of a “normal” CV, i.e. there is no spectroscopic signature related to the nova eruption left.
@@
This system is also known as Gaia17cbl. The two SDSS spectra show double-peaked Balmer emission lines. Whereas the spectrum taken on $\mathrm{MJD}=59709$ was obtained during quiescence, the one taken one spectrum ($\mathrm{MJD}=59737$) shows a continuum flux level higher by a factor of about eight, and also shows a strong double-peaked He ii emission line. The ZTF and Gaia light curves show frequent outbursts ($\Delta m\simeq 2$). This is a dwarf nova, and the second SDSS spectrum was obtained during one of the outbursts.
@@
This is the SU UMa dwarf nova V550 Cyg. The SDSS spectrum shows a red continuum with narrow Balmer absorption lines. The spectrum is consistent with the photometric SED and the synthetic magnitude is consistent with the Gaia photometry. The object is located within the main sequence of the HR diagram when reddening ($0.69$) is accounted for. This is all inconsistent with the properties of an SU UMa dwarf nova. CXOGSG J200505.2+322121 is an X-ray source detected 3.8 arcsec away from the catalogue position of V550 Cyg, and we tentatively identify Gaia EDR3 2031052139821246336 as the counterpart to this SU UMa dwarf nova, with a parallax of 2.83 mas and a quiescent magnitude of $G\simeq 20.8$. It is likely that V550 Cyg is actually Gaia EDR3 2031052139821246336  – Skiff (1999) refers to it being the “southeastern of a pair” and it is in a crowded area close to the Galactic plane.
@@
This system is also known as ZTF J200729.55+625047.6. The SDSS spectrum shows asymmetric Balmer and He i emission lines indicative of an accreting magnetic white dwarf. The spectrum and SED are dominated by the donor star. The object is located near the main sequence. The ZTF light curve (Fig. LABEL:fig:J2007+6250) shows frequent state changes, which are consistent with our tentative classification of J2007+6250 as a magnetic CV. We obtained a period of $5.61(2)$ h from the TESS light curve which we believe is more reliable than the $72$ d period quoted by Chen et al. (2020).
@@ 6.1 Effectiveness of the different CV targeting strategies within SDSS-V
In Section 3 the use of SDSS-V cartons was described noting that cartons can overlap in terms of their target contents. Fig. LABEL:fig:UpSet illustrates the relative contribution of the targeting methods as implemented in the SDSS cartons that resulted in the identification of CVs. The left-most five histograms show the number of CVs that were identified only thanks to a single carton, i.e. they were not part of any overlap between the different cartons. The other histograms to the right show the number of CVs that were targeted by multiple cartons, e.g. the right-most histogram illustrates that only a single CV was targeted by all cartons. It is interesting to note that, as expected, targeting known or candidate CVs, mwm_cb_cvcandidates, naturally accounted for the largest number (364) of CVs observed by SDSS-V, a large fraction of those ($\simeq 44$ per cent) were not contained within any other carton. Inspection of these systems shows that this was because of the lack of ultraviolet photometry and/or poor Gaia astrometry. The various ultraviolet-excess target selections (mwm_cb_gaiagalex and mwm_cb_uvex1-5) were also very successful in leading to the identification of CVs. Whereas the carton mwm_wd_core principally targets single white dwarfs selected from the Gaia data (Gentile Fusillo et al., 2019, 2021), it contains a number of low-accretion rate CVs with little to no spectroscopic signature of the donor star, primarily WZ Sge dwarf novae and AM CVn systems. Twenty CVs were observed by SDSS-V solely because of their properties in the eROSITA survey with all but two arising from the BHM_SPIDERS_AGN cartons. These twenty were predominantly faint systems ($m>19$) which explains why eight did not have Gaia counterparts and fourteen could not be completely classified. Our eROSITA-related observations were all obtained based on an early release of eROSITA data (eFEDS) with limited coverage and the proportion of eROSITA-related CV discoveries is expected to rise following the larger eRASS1 release. Finally, only eight CVs are entirely serendipitous discoveries. All eight were targeted by “open fibre cartons”, different projects making use of spare fibres left over once all main survey targets are allocated in a given field. The specific open fibre programs in this case were a representative spectroscopic mapping of the Gaia HR diagram and searches for quasars as well as for detached, spatially unresolved white dwarf plus main sequence star binaries.
@@ 6.2.2 Specifically targeting CVs
In SDSS I to IV, CVs were not specifically targeted and discoveries were essentially random. Inight et al. (2023a) therefore assessed, and corrected for, the incompleteness of the CVs in SDSS I to IV by considering the proportion of known CVs from Ritter & Kolb (2003) that were rediscovered in SDSS I to IV. In SDSS-V we take a different approach. Fig. LABEL:fig:UpSet shows that of the SDSS-V 505 CVs all but the eight “other” CVs were contained in one or more of the CV-related cartons from which we calculate that $98.4$ per  cent of the CVs in the footprint are targeted. Completeness is therefore determined by aggregating the targets in the CV-related cartons and calculating the proportion that have actually been observed.
@@ 6.2.3 Space density estimates
We followed the approach in Section 6.4 in Inight et al. (2023a) using the subset of CVs from this work with $|b|>20\degree$ to avoid the Galactic plane. A footprint was established consisting of the sky coverage (defined in terms of HEALPixs) of all the fields with $|b|>20\degree$ containing at least one observation in a CV-related carton. This footprint took account of the different diameters in the focal plane of plates ($1.49\degree$) and fields at Apache Point observatory ($1.414\degree$) and Los Campanos observatory ($0.9465\degree$). Each carton was then analysed to find the total number of targets within the footprint and the proportion that had been observed and hence the overall completeness (see Table LABEL:tab:completeness). The assumptions from Inight et al. (2023a) regarding scale heights and brightness were used to compute effective volumes and hence space densities. The results are tabulated in Table LABEL:tab:spacedensity and plotted in Fig. LABEL:fig:space_densities. They are comparable with, and reinforce, previous findings (Warner, 1995; Pretorius et al., 2007b; Pretorius & Knigge, 2012; Pretorius et al., 2013; Schwope, 2018; Pala et al., 2020).
@@ 7 Conclusions
We have used a convolutional neural network (CNN) to analyse the complete set of spectroscopic data from SDSS-V resulting in the discovery of 62 new CVs, spectroscopic confirmation of 243 CV candidates and 68  new or improved orbital periods. We have demonstrated that the CNN recovers at least 90 per cent of CVs within a sample whilst rejecting at least 98 per cent of the non-CVs. Furthermore we have shown that, with the exception of pre-polars, there is no significant selection bias between CV sub-types. This approach can be refined, using our enlarged sample of known CVS for training, and used not only for future SDSS-V surveys but potentially also for other spectroscopic surveys such as DESI, WEAVE and 4MOST.
@@ 8 Acknowledgements
This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 101020057). MRS acknowledges support from Fondecyt (grant 1221059).
@@ A.1 Model assessment
An illustration of Receiver Operating Characteristic (ROC) curves showing the ideal curve (purple) and a typical curve (blue). The green dotted curve represents an untrained CNN that is effectively making random predictions. The Area Under the Curve (AUC) for the typical case is shaded light blue.
@@ A.2.1 The number of convolutional and dense layers
Heat map showing the effect of increasing the number of kernels and the width of each kernel. The stochastic variation in the output of the CNN model is very evident even though each estimate was repeated five times and the median taken. The choice of three kernels and a width of 25 results is a reasonable compromise between quality and overfitting.

### v2

@@ 4.3.3 CV sub-types and potential selection bias
Thirdly we considered the BOSS QA sample. This is essentially an unseen sample and shows very little bias. The overall recovery rate for the BOSS QA sample is 90 per cent.
@@
ID Sp $P$
J000155.14–670743.3 EF Tuc 4707482485122305536 3.5: 14.85 $1329$ NL 14 13 24
J000600.15+012129.8 2738755406045571968 1.52(10) 19.57 $352$ WZ Sge: 15 15 15
J000633.70–690033.8 HV 8001 4706297108508032128 1.90(3) 15.1 $241$ SU UMa 9 17 17
J000844.33–014014.8 Gaia15abi 2544817841421994624 18.15 $1236$ NL:/Polar 2 26
J001204.50+020129.6 2546760133008036096 3.288(1) 16.45 $1667$ NL 21 18 1
J002243.55+061002.7 MGAB-V295 2747622177049417344 1.9(1): 19.61 $2113$ SU UMa 30 26 26
J002257.65+614107.5 V1033 Cas 430446643235641472 4.033 16.58 $1557$ IP 5 10 10
J002728.02–010828.7 EN Cet 2541910801397761152 1.424(1) 20.76 $1218$ WZ Sge 12 27 6
J002848.83+591722.0 V709 Cas 428101969045877504 5.33290(7) 14.39 $725$ IP 8 8 11
@@
J014227.07+001729.8 2510205490257050496 1.88(4) 19.89 $697$ Polar: 22 22
This is a small sample of the data. The full table can be found in the supplementary data. References:  1 This work, 2 Hodgkin et al. (2021a), 3 Szkody et al. (2014), 4 Southworth et al. (2010), 5 Halpern & Mirabal (2006), 6 Dillon et al. (2008), 7 AAVSO, 8 Motch et al. (1996), 9 Luyten (1933), 10 Bonnet-Bidaud et al. (2007), 11 Thorstensen et al. (2010), 12 Bao Supernova Survey et al. (1997), 13 Stobie et al. (1995), 14 Chen et al. (2001), 15 Inight et al. (2023b), 16 VSNET 12318, 17 Augusteijn et al. (2010), 18 Luo et al. (2018), 19 Szkody et al. (2004), 20 Drake et al. (2009b), 21 vsnet-chat 7946, 22 Carbon et al. (2017), 23 Szkody et al. (2018), 24 CBA512, 25 Stefanov & Stefanov (2023), 26 Inight et al. (2023a), 27 Szkody et al. (2005), 28 Kato et al. (2015), 29 Southworth et al. (2007), 30 MGAB Variable Star Catalog, 31 Southworth et al. (2008), 32 Shappee et al. (2014)
@@ 5 CVs found in SDSS-V
The final collection of CVs observed by SDSS-V consists of the 422 from Section 4.1, the 55 from Section 4.2, 21 CVs from the BOSS QA sample that were not found by the CNN and 6 CVs from mwm_cb_cvcandidates that were also not found by the CNN. These 504  CVs were further analysed following the methods in Inight et al. (2023a, b). This analysis of the SDSS-Vspectroscopy and ancillary data of the 504  CVs in our sample yielded 61  new CVs, spectroscopic confirmation of 248 candidates and 82  new or improved orbital periods. Prior to SDSS-V, one of the authors (JRT) obtained time-resolved spectroscopy using the MDM observatory at Kitt Peak (see Thorstensen 2020 for more details on the instrument and data reduction techniques) of a number of our CVs. We present here the so far unpublished periods resulting from these observations. The full set of data is included in the supplementary data and a sample is shown in Table LABEL:tab:cvlist. Whilst analysing the 504  CVs in SDSS-V some CVs with unusual or noteworthy characteristics were identified and these are discussed below. For simplicity the names of CVs have been abbreviated in the text by dropping the arcseconds; so for example SDSS J015543.47+002807.4 becomes J0155+0028.
@@
The CRTS light curve (Fig. LABEL:fig:J0155+0028) captured the system mainly in a high state ($16-18$ mag), in contrast, the more recent ZTF data shows the system predominantly in a low state ($\simeq 20$ mag), with two relatively short high states. The transitions between high and low states can occur on time scales of days,
@@
This system is also known as ATO J119.2035--12.7816. The SDSS spectrum shows double-peaked Balmer and He i emission lines. The object is on the main sequence in the HR diagram just below the region occupied by G-type stars. GALEX observations show that the system is fairly bright in the ultraviolet. The ZTF light curve exhibits outbursts. We obtained a period of $10.79(4)$ h from the TESS light curve. This appears to be an U Gem dwarf nova, possibly with an evolved donor although an unevolved G-type donor would also be consistent.
@@
This system is also known as BMAM-V636. The SDSS spectrum shows Balmer, He i and He ii emission lines. J1618+0309 is a known polar (BMAM-V636). Cyclotron humps are evident in the spectrum (Fig. LABEL:fig:J1618+0309) indicating $B\simeq 37$ MG.
@@
This is the classical nova V728 Sco, which erupted in 1862. The SDSS spectrum shows double-peaked Balmer, He i and He ii emission lines. J1739–4527 is located close to the main sequence in the HR diagram. There is only ATLAS light curve data available. We concur with Tappert et al. (2012) that J1739–4527 now exhibits the characteristics of a “normal” CV, i.e. there is no spectroscopic signature related to the nova eruption left.
@@
This system is also known as Gaia17cbl. The two SDSS spectra show double-peaked Balmer emission lines. Whereas the spectrum taken on $\mathrm{MJD}=59709$ was obtained during quiescence, the one taken one spectrum ($\mathrm{MJD}=59737$) shows a continuum flux level higher by a factor of about eight, and also shows a strong double-peaked He ii emission line. The ZTF and Gaia light curves show frequent outbursts ($\Delta m\simeq 2$). This is a U Gem, and the second SDSS spectrum was obtained during one of the outbursts.
@@
This is the SU UMa dwarf nova V550 Cyg. The SDSS spectrum shows a red continuum with narrow Balmer absorption lines. The spectrum is consistent with the photometric SED and the synthetic magnitude is consistent with the Gaia photometry. The object is located within the main sequence of the HR diagram when reddening of ($E(B-V)=0.69$) is accounted for. This is all inconsistent with the properties of an SU UMa dwarf nova. CXOGSG J200505.2+322121has a textitChandra X-ray detection in the range $0.1-10$ keV at 3.8 arcsec away from the catalogue position of V550 Cyg, and we tentatively identify Gaia EDR3 2031052139821246336 as the counterpart to this SU UMa dwarf nova, with a parallax of 2.83 mas and a quiescent magnitude of $G\simeq 20.8$. It is likely that V550 Cyg is actually Gaia EDR3 2031052139821246336  – Skiff (1999) refers to it being the “southeastern of a pair” and it is in a crowded area close to the Galactic plane.
@@
This system is also known as ZTF J200729.55+625047.6. The SDSS spectrum shows Balmer and He i emission lines indicative of an accreting magnetic white dwarf. The spectrum and SED are dominated by the donor star. The object is located near the main sequence. The ZTF light curve (Fig. LABEL:fig:J2007+6250) shows frequent state changes but no apparent outbursts, which are consistent with our tentative classification of J2007+6250 as a magnetic CV. We obtained a period of $5.610679(1)$ h from the TESS light curve which we believe is more reliable than the $72$ d period quoted by Chen et al. (2020).
@@ 6.1 Effectiveness of the different CV targeting strategies within SDSS-V
In Section 3 the use of SDSS-V cartons was described noting that cartons can overlap in terms of their target contents. Figure LABEL:fig:UpSet illustrates the relative contribution of the targeting methods as implemented in the SDSS cartons that resulted in the identification of the 504  CVs. The left-most five histograms show the number of CVs that were identified only thanks to a single carton, i.e. they were not part of any overlap between the different cartons. The other histograms to the right show the number of CVs that were targeted by multiple cartons, e.g. the right-most histogram illustrates that only a single CV was targeted by all cartons. It is interesting to note that, as expected, targeting known or candidate CVs, mwm_cb_cvcandidates, naturally accounted for the largest number (364) of CVs observed by SDSS-V, a large fraction of those ($\simeq 44$ per cent) were not contained within any other carton. Inspection of these systems shows that this was because of the lack of ultraviolet photometry and/or poor Gaia astrometry. The various ultraviolet-excess target selections (mwm_cb_gaiagalex and mwm_cb_uvex1-5) were also very successful in leading to the identification of CVs. Whereas the carton mwm_wd_core principally targets single white dwarfs selected from the Gaia data (Gentile Fusillo et al., 2019, 2021), it contains a number of low-accretion rate CVs with little to no spectroscopic signature of the donor star, primarily WZ Sge dwarf novae and AM CVn systems. Twenty CVs were observed by SDSS-V solely because of their properties in the eROSITA survey with all but two arising from the BHM_SPIDERS_AGN cartons. These twenty were predominantly faint systems ($m>19$) which explains why eight did not have Gaia counterparts and fourteen could not be completely classified. Our eROSITA-related observations were all obtained based on an early release of eROSITA data (eFEDS) with limited coverage and the proportion of eROSITA-related CV discoveries is expected to rise following the larger eRASS1 release. Finally, only eight CVs are entirely serendipitous discoveries. All eight were targeted by “open fibre cartons”, different projects making use of spare fibres left over once all main survey targets are allocated in a given field. The specific open fibre programs in this case were a representative spectroscopic mapping of the Gaia HR diagram and searches for quasars as well as for detached, spatially unresolved white dwarf plus main sequence star binaries.
@@ 6.2.2 Specifically targeting CVs
In SDSS I to IV, CVs were not specifically targeted and discoveries were essentially random. Inight et al. (2023a) therefore assessed, and corrected for, the incompleteness of the CVs in SDSS I to IV by considering the proportion of known CVs from Ritter & Kolb (2003) that were rediscovered in SDSS I to IV. In SDSS-V we take a different approach. Figure LABEL:fig:UpSet shows that of the SDSS-V 504 CVs all but the eight “other” CVs were contained in one or more of the CV-related cartons from which we calculate that $98.4$ per  cent of the CVs in the footprint are targeted. Completeness is therefore determined by aggregating the targets in the CV-related cartons and calculating the proportion that have actually been observed.
@@ 6.2.3 Space density estimates
We followed the approach in Section 6.4 in Inight et al. (2023a) using the subset of CVs from this work with $|b|>20\degree$ to avoid the Galactic plane. A footprint was established consisting of the sky coverage (defined in terms of HEALPixs) of all the fields with $|b|>20\degree$ containing at least one observation in a CV-related carton. This footprint took account of the different diameters in the focal plane of plates ($1.49\degree$) and fields at Apache Point observatory ($1.414\degree$) and Los Campanos observatory ($0.9465\degree$). Each carton was then analysed to find the total number of targets within the footprint and the proportion that had been observed and hence the overall completeness (see Table LABEL:tab:completeness). The assumptions from Inight et al. (2023a) regarding scale heights and brightness were used to compute effective volumes and hence space densities. The uncertainty in the estimate for space density is dominated by the assumptions regarding scale height and we follow the approach used in Inight et al. (2023a) to provide a notional error estimate using different scale height assumptions. The results are tabulated in Table LABEL:tab:spacedensity and plotted in Fig. LABEL:fig:space_densities. They are comparable with, and reinforce, previous findings (Warner, 1995; Pretorius et al., 2007b; Pretorius & Knigge, 2012; Pretorius et al., 2013; Schwope, 2018; Pala et al., 2020).
@@ 7 Conclusions
We have used a convolutional neural network (CNN) to analyse the complete set of spectroscopic data from SDSS-V resulting in the discovery of 61 new CVs, spectroscopic confirmation of 248 CV candidates and 82  new or improved orbital periods. We have demonstrated that the CNN recovers at least 90 per cent of CVs within a sample whilst rejecting at least 98 per cent of the non-CVs. Furthermore we have shown that, with the exception of pre-polars, there is no significant selection bias between CV sub-types. This approach can be refined, using our enlarged sample of known CVS for training, and used not only for future SDSS-V surveys but potentially also for other spectroscopic surveys such as DESI, WEAVE and 4MOST.
@@ 8 Acknowledgements
This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 101020057). MRS acknowledges support from Fondecyt (grant 1221059). G.T. was supported by the IN109723 grant from the Programa de Apoyo a Proyectos de Investigación e Innovación Tecnológica (PAPIIT).